### v1

# Epistemic diversity across language models mitigates knowledge collapse

Abstract.
The growing use of artificial intelligence (AI) raises concerns of knowledge collapse, i.e., a reduction to the most dominant and central set of ideas. Prior work has demonstrated single-model collapse, defined as performance decay in an AI model trained on its own output. Inspired by ecology, we ask whether AI ecosystem diversity, that is, diversity among models, can mitigate such a collapse. We build on the single-model approach but focus on ecosystems of models trained on their collective output. To study the effect of diversity on model performance, we segment the training data across language models and evaluate the resulting ecosystems over ten, self-training iterations. We find that increased epistemic diversity mitigates collapse, but, interestingly, only up to an optimal level. Our results suggest that an ecosystem containing only a few diverse models fails to express the rich mixture of the full, true distribution, resulting in rapid performance decay. Yet distributing the data across too many models reduces each model’s approximation capacity on the true distribution, leading to poor performance already in the first iteration step. In the context of AI monoculture, our results suggest the need to monitor diversity across AI systems and to develop policies that incentivize more domain- and community-specific models.
artificial intelligence, epistemic diversity, model collapse, AI ecosystem, system pluralism, distributed artificial intelligence
This work is supported by the University of Washington’s Center for an Informed Public and the John S. and James L. Knight Foundation (G-2019-58788).
1. Introduction
Figure 1. Illustration of model collapse in AI ecosystems. Because of a model’s (m) inherent biases (in learning and resampling) and finite data, it cannot perfectly regenerate the information space (e.g., concepts, claims, facts) it was trained on. Some information becomes distorted, and some is lost entirely due to low sampling probabilities. When this training–generation cycle is repeated, the model’s output becomes increasingly homogeneous and degrades over iterations $t$. The probability distribution of the information space expressed in the output (and represented by the model, $p_{\theta}$) gradually narrows with increasing $t$, converging toward delta-like functions. Collapse can occur at both the individual model level and the ecosystem level, meaning that even a collection of models trained on their collective output can eventually degenerate (here illustrated with three models). In society, we expect knowledge to grow and improve. In contrast, without human intervention, the best we can hope for from AI models is that they preserve the distribution of information they are originally trained on ($p_{data}$). Thus, for the purpose of this paper, we evaluate ecosystems based on model performance on the original dataset, averaged over 10 training–generation iterations.
AI’s growing role in knowledge production and curation raises concerns about a potential collapse of human knowledge (Peterson, 2025; Sourati et al., 2025; Wagner and Jiang, 2025). The scenario unfolds as follows: as AI systems become widely incorporated in private and professional contexts (e.g., in scientific discovery and writing (Liang et al., 2025; Mishra et al., 2024; Messeri and Crockett, 2024)), they also generate an increasing volume of artificial data used for training. This amplifies the risk of model collapse, a degenerative process in which generative models re-trained on their own outputs begin producing homogeneous, biased, and nonsensical outputs (Herel and Mikolov, 2024; Shumailov et al., 2024; Bertrand et al., 2024). Ultimately, such homogenization in model-generated data, particularly with language, may contribute to knowledge collapse, a reduction of human knowledge to the most dominant and central set of ideas (Peterson, 2025; Wright et al., 2025).
Even before AI, philosophers have argued for the benefits of epistemic diversity in science, not only for advancing knowledge production but also to prevent injustices such as the undervaluing of minority perspectives (Douglas, 2009; Longino, 1990; Wright et al., 2025; Rolin, 2019; Solomon, 2006). When scientists learn from both their own results and one another’s findings, diversity reduces the risk of converging on false beliefs. For example, the myth that Sickle Cell Disease is specific to Black Americans persisted in medicine for decades (Anderson et al., 2023). Simulations of collective knowledge production show that diversity is beneficial, not monotonically, but up to a certain level; if diversity increases to the point where scientists no longer learn from each other, the overall success rate begins to decline as well [^1] To clarify our position: We do not believe this upper bound has ever been reached. (Zollman, 2010; Wu and O’Connor, 2023; Fazelpour and Steel, 2022).
Like scientists exchanging new findings, AI models learn from their own output and the output of other AI [^2] We are not attempting to compare humans and AI models underlying learning mechanisms but only to illustrate information interactions in human and AI ecosystems. We use terms like “learning” and “training” because they are established in AI research.. In the presence of AI monoculture—where a few large models, trained on similar datasets with similar model architectures, (Lu and Kleek, 2024; Lazar and Nelson, 2023; Sourati et al., 2025; Bommasani et al., 2022) dominate the AI ecosystem—we ask whether such an optimal degree of diversity exists in AI ecosystems that could mitigate monocultural shifts.
In homogeneous ecosystems, model collapse occurs because of a model’s non-zero probability of losing information or learning wrong patterns, which is irreversible, when relying only on its own output (Shumailov et al., 2024). Analogous to diversity in human systems, we hypothesize that diverse models trained on their collective outputs can benefit their performance by increasing distributional variance in artificial data generation. Viewing an AI model as an epistemological representation, we use the Hill-Shannon Diversity (HSD, $D$) (Roswell et al., 2021; Jost, 2006; Hill, 1973) to measure ecosystem diversity, which quantifies the effective number of equally frequent, diverse AI models. It is important to note that we manipulate only the diversity across models, without altering the diversity of the dataset on which the models are initially trained, as has been studied in prior work (e.g., Wright et al., 2025; Schaffelder and Gatt, 2025; Zhu et al., 2025).
To address the challenges of studying the evolution of the entire AI ecosystem (e.g., ChatGPT, Gemini, Claude, etc., and all their inputs and outputs), we build upon previous work that analyzes a single-model collapse (e.g. Shumailov et al., 2024; Herel and Mikolov, 2024). This previous work re-trains small language models on their own output over multiple iterations and measures performance using perplexity. We take a similar approach, but instead focus on a collection of models, or a model ecosystem, rather than a single model. This then allows us to study the effects of our main variable of interest—ecosystem diversity—on model performance. We measure this effect separately on two different model types (OPT-125m (Zhang et al., 2022) and GPT2 (Radford et al., 2019)).
While there are many forms of diversity across models (e.g., size, architecture, alignment, system messaging) [^3] A specific model architecture can be used to build different AI systems (i.e., the end products with which users interact). For simplicity, we use “model type” to differentiate between OPT-125m and GPT2 pre-trained base models and “model” to refer to the fine-tuned versions., we focus on training data as the way of instantiating epistemic diversity. To do this under a controlled setting, we segment the training data, which is fixed in size and, in the first iteration step, also in content. Through this process, we go from one, pre-trained model to multiple fine-tuned models (MM=1, 2, 4, 16), as shown in Figure 2. We use Wikitext2 (Merity et al., 2016) as the dataset for generating artificial training data, evaluation, and initial training. This segmentation approach allows us to vary ecosystem diversity ($D=M$) in a relatively simple manner, while keeping factors, such as model type and the per-ecosystem resource budget for sample size and computation, fixed.
Our experiments under a fixed budget yield an optimal ecosystem diversity greater than zero for model performance when models are trained on their own output, as shown in Figure 3. Too much diversity (and less training data per model) results in high perplexity at early stages, and too little leads to a high degeneration rate in the long term. We theorize that a trade-off primarily between increasing ecosystem-level expressivity (the ensemble can represent a richer mixture as members can specialize to subpopulations) and decreasing approximation capacity of the individual training data on the true distribution explains the U-shaped relationship between performance and diversity.
The optimal ecosystem diversity observed in our experiments suggests that a similar optimum may exist in reality. As AI models become increasingly incorporated into major institutions such as science, education, government, and economics, a collapse toward the most dominant ideas could have severe societal implications. These risks extend beyond knowledge production and may also impact domains of epistemic and social justice.
2. Background and related work
2.1. Collapse
Model collapse has been theoretically and/or empirically demonstrated for generative models in various settings (Alemohammad et al., 2023; Shumailov et al., 2024; Herel and Mikolov, 2024; Zhang et al., 2025b; Xing et al., 2025; Dohmatob et al., 2024b). The observed effects typically involve some form of performance decay; however, the field has not yet arrived at a general definition of the phenomenon (Schaeffer et al., 2025). Our experimental approach for the single model follows Shumailov et al. (2024), who first formalized model collapse and revealed it in OPT-125m. Model collapse arises primarily from three sources of error related to training data quality (precision and recall), the model’s capacity to represent a given distribution (to “express” it), and the model’s capacity to accurately learn what it is capable of representing (Shumailov et al., 2024): statistical approximation error, functional expressivity error, and functional approximation error. We will return to these errors when discussing differences in model performance across collections of diverse models. Model collapse is not only a threat to generative models accidentally (or inevitably) trained on artificial internet data but also when considering approaches leveraging synthetic data for efficiency reasons (Bai et al., 2022; Goyal and Mahmoud, 2024; Feng et al., 2024a). Very recent work theorizes that model collapse will result in a collapse of knowledge, defined as the continuous narrowing down of ideas, which will lead to broader epistemic problems for our society (Wright et al., 2025; Peterson, 2025; Wagner and Jiang, 2025).
2.2. Mitigation and prevention
Because of the epistemic risks of collapse, scholars have developed various strategies to prevent such performance decay, or, put differently, to ensure stability.
The majority of intervention strategies focus on data curation, primarily at the data level, including adding fresh data and filtering or correcting artificial data (Gerstgrasser et al., 2024; Bertrand et al., 2024; Amin et al., 2025; Xing et al., 2025; Fu et al., 2025; Zhu et al., 2025; Dohmatob et al., 2024b; Feng et al., 2024a). While such methods have been shown to mitigate the problem, in most experimental settings they cannot entirely prevent performance decay, even when the involved technical and practical challenges of constantly sourcing fresh data or reliably separating real from artificial data set aside (Xing et al., 2025; Dohmatob et al., 2024a).
2.3. AI ecosystem
Our real AI ecosystem does not consist of a single self-training model but rather a collection of models trained on both their own outputs and those produced by others. Thus, more closely aligned with the work presented here are recent approaches that seek to understand model collapse at the ecosystem level––for instance, by simulating interactions among AI models via retrieval-augmented generation (RAG) (Wang et al., 2025), or by sourcing artificial data from multiple language models for training and evaluating a single model (Schaffelder and Gatt, 2025). In contrast, our approach focuses on diversity across models as the independent variable, involving real (non-simulated) re-training of language models on their collective output.
2.4. Diversity
Epistemic diversity in humans refers to different ways of knowing. It broadly describes the phenomenon that individuals with different background assumptions, shaped by diverse lived experiences, values, and beliefs, can arrive at different interpretations when presented with the same information or data (Douglas, 2009; Longino, 1990; Solomon, 2006).
AI models generate information according to learned probability distributions, which means each model reflects specific assumptions about what counts as relevant, likely, or meaningful knowledge; in other words, a particular epistemology. Epistemic diversity in an AI ecosystem therefore captures the extent to which multiple models, shaped by different data sources, architectures, or objectives, may yield divergent interpretations or outputs when presented with the same input. In recent AI research, epistemic diversity within and across AI models has attracted increased interest. For example, scholars have shown that diverse models collectively make better decisions (Kleinberg and Raghavan, 2021), reduce the risk of failures (Lu and Kleek, 2024), and enable fairer representations of diverse societies (Fisher et al., 2025; Varon et al., 2024; Gebru and Torres, 2024; Jain et al., 2024; Blili-Hamelin et al., 2025).
Originating in ecology, a common method to measure such diversity is the Hill–Shannon Diversity (HSD) (Roswell et al., 2021; Jost, 2006; Hill, 1973; Wright et al., 2025), which reflects the effective number of diverse, equally common elements (e.g., models, species, or categories) in a population.
Figure 2. Instantiating ecosystem diversity and evaluating model performance. We examine the performance of models $m$ within collections of increasing diversity $D$ when trained on collective output across 10 training iterations $t$. The number of models $M$ equals the Hill-Shannon Diversity $D$, which is instantiated through segmentation of the collected output. At each iteration, we concatenate and shuffle the $M$ generated datasets, and uniformly redistribute to the $M$ models for the next step. This segmentation approach keeps the number of tokens per ecosystem NN fixed, yielding a per-model sample size of $n=N/M$. The ground-truth dataset used to measure performance, the original Wikitext2 test split, is fixed across all models and iterations.
3. Approach
We study the progress of model performance in AI ecosystems containing 1, 2, 4, and 16 models, each re-trained on their collective output over 10 iterations. All models begin as identical copies of a pre-trained model but are fine-tuned on unique subsets of the training data. This segmentation approach enables us to isolate and examine the effect of diversity across models on their performance under fixed model architecture and a fixed per-ecosystem resource budget for data size (number of tokens, $N$) and compute.
Each iteration step involves an evolving training dataset used to update model weights, a fixed generation dataset used to generate artificial training data for the next iteration, and a fixed evaluation dataset used to measure model perplexity (see Figure 2). The evaluation dataset is the same for all models and corresponds to the original Wikitext2 test set. The generation datasets are unique to each model but fixed across all iterations, corresponding to randomly sampled subsets of the original Wikitext2 training data.
In the first iteration step, the training data is identical to the generation dataset for a given model. Thereafter, it evolves in content because it is sourced from the outputs of the previous iteration: At each iteration, the $M$ generated datasets are concatenated, shuffled, and uniformly redistributed to the $M$ models for the next round. Throughout the ten iterations, the training and generation datasets for each model within an ecosystem are equal in size, $n=N/M$. For example, for two models ($D=M=2$), the sizes of the training and generation datasets are each half of the original size, $n=N/2$ (see Figure 2). In the case of the single-model ecosystem ($M=1$), our approach parallels the methods used in prior work on single-model collapse (e.g., Shumailov et al., 2024).
Translated to the real world, the segmentation approach broadly mirrors domain- or community-specific models (e.g., Feng et al., 2024b; Tanksley et al., 2025) trained on distinct segments of internet data. While tailored to different communities that might use different inputs, the generated outputs are collectively stored in a single data pool, from which future training samples are drawn.
3.1. Ecosystem diversity $D$ and model performance
Drawing from ecology, we use the HSD to measure ecosystem diversity, that is, diversity across models. The HSD quantifies the effective number of equally common elements (models, species, or categories) in a population (Roswell et al., 2021; Jost, 2006; Hill, 1973). Let $M$ denote the number of models, then we calculate HSD as:
(1) $D(M)=\exp\!\left(-\sum_{m=1}^{M}w_{m}\ln w_{m}\right)$
Because of equal weight across models ($w_{m}=1/M$), the diversity simplifies to the number of models in a given ecosystem:
(2) $D(M)=M$
Consistent with prior work (Shumailov et al., 2024), we compute each model’s perplexity on the original Wikitext-2 test split to assess model performance. Perplexity reflects a model’s ability to complete the test set text sequences (lower is better) and indicates the degree of alignment between the model’s probability distribution $p_{\theta}$ and the true distribution $p_{data}$. The width of the perplexity distribution serves as an indicator of model collections’ expressivity, which we use in Figure 5.
In an AI ecosystem composed of $M$ different models, we calculate the combined mean perplexity $\mu$ for a given iteration step $t$ as:
(3) $\mu_{t}=\frac{1}{M}\sum_{m=1}^{M}\mu_{m,t}(N_{test})\text{ , }$
whereas $\mu_{m,t}(N_{test})$ corresponds to the mean perplexity across the text sequences of the original Wikitext2 test split.
To compare ecosystems across iterations, we calculate an aggregated mean perplexity. Let $T$ denote the number of iteration steps, $\mu_{t}$ an ecosystem’s mean perplexity for a given iteration step $t$, then, the aggregated perplexity $\mu_{T}$ equals:
(4) $\mu_{T}=\frac{1}{T}\sum_{t=1}^{T}\mu_{t}$
3.2. Generation and training details
We use GPT2 (Radford et al., 2019) and OPT-125m (Zhang et al., 2022), two open-accessible, small, and well-studied language models used in prior work on collapse (e.g., Shumailov et al., 2024; Herel and Mikolov, 2024).
Analogous to prior work (Herel and Mikolov, 2024; Shumailov et al., 2024), we use Wikitext2 (Merity et al., 2016) as the original dataset. To reduce computational effort during training and generation, we use a randomly sampled subset of two-fifths of the original size (=$N$), while keeping the original Wikitext2 validation and test splits for evaluation. Using a block size of 128 tokens, the sizes of the training–generation, validation, and test splits correspond to 7,500, 1,900, and 2,200 data points, respectively.
Our approach for generating the artificial dataset follows a method introduced in prior work (Shumailov et al., 2024). For each available token sequence of length 128 in the generation dataset, we prompt the model to generate a continuation sequence of equal length, using a five-way beam search. Because we iterate through the entire generation dataset, the model would produce the original generation dataset if it had zero error. At each iteration, each model is trained for 5 epochs with a learning rate of $1e^{-5}$ on a training sample of size $n=N/M$, using a block size of 128 tokens. The best-performing model on the validation set is preserved for subsequent iterations.
All experiments are carried out using the HuggingFace Transformers library (Wolf et al., 2019) and the Python Lightning framework (Falcon and The PyTorch Lightning team, 2019).
4. Results
We conducted various experiments, testing the impact of increasing diversity on ecosystems of AI models trained on their collective output. Under a fixed data size and compute budget, we find that performance improves for increased diversity up to a certain point. In other words, we find an optimal degree of diversity greater than zero (see Figure 3).
Figure 4 shows the mean perplexity for the OPT-125m and GPT2 collections of 1, 2, 4, and 16 models trained on unique segments of the artificial training data across 10 re-training iterations. Ecosystems with a diversity of four ($D=M=4$) achieve the lowest aggregated mean perplexity for both model types, with the effect more pronounced for OPT-125m. Ecosystems with lower diversity ($M=1,2$) perform well in early stages but show rapid perplexity increases in later iterations.
The least diverse ecosystem ($D=1$) is the best performer only at $t=0$, when all models are trained on real data [^4] Progress, including perplexity for the original models, is shown in Figure 7 in the appendix.. The only difference is the per-model sample size ($n=N/M$). Conversely, the ecosystem with the highest diversity ($M=16$) begins with relatively high perplexity, but, after evolving, it approaches similarly low perplexity values as $M=4$. In addition to aggregated mean perplexity, diversity also affects the rate of perplexity change after 10 iterations. While low-diversity ecosystems exhibit a clearly positive rate, indicating a strong performance decay, the highest-diversity ecosystem yields a negative rate (see Figure 4).
Figure 5 shows the mixed perplexity distributions of the GPT2 and OPT-125m ecosystems on the original training dataset, indicating that collections with greater diversity exhibit both greater variety in the information represented and lower alignment with the true distribution.
Figure 3. AI ecosystems exhibit an optimal degree of diversity (Hill–Shannon Diversity $D$) for performance, operationalized as the mean perplexity on the original test set averaged across 10 training–generation iterations. For both model types, GPT2 and OPT-125m, this perplexity minimum occurs at $D=M=4$. The colored symbols correspond to the colors in Figure 2, which indicate different levels of diversity (e.g., orange corresponds to $D=2$).
(a) OPT-125m
(b) GPT2
Figure 4. Diversity across models $D$ mitigate model collapse. Comparison of ecosystems’ perplexity with increasing diversity ($D=M$) at fixed dataset size across ten iteration steps. For both model types, $M=4$ yields the lowest mean perplexity at the early and late stages of evolution. After 10 iterations, the perplexity increasing rate is positive for low diversity ($M=1,2$), near 0 for $m=4$, and negative for the highest diversity ecosystem ($M=16$). The shaded region is a combined standard deviation across the models and data points.
(a) OPT-125m
(b) GPT2
Figure 5. Perplexity distributions widen with increasing diversity. Mixed perplexity distribution on the original Wikitext2 train data points, evaluated by all models in a given ecosystem with diversity ($D$ = $M$ = 1, 2, 4, 16) at iteration steps $t$ = 0 (left) and $t$ = 9 (right). As ecosystems’ diversity increases, the distributions become wider indicating greater expressivity. Simultaneously, their means shift toward higher perplexity, suggesting greater divergence from the distribution in the original training set. Across all ecosystems, the distributions narrow from the beginning to the end of the model iterations.
5. Theoretical interpretation
The inverted U-shaped relationship between diversity and performance reflects a trade-off between opposing effects of diversity across models. To understand these effects, we draw on the three sources of error in model collapse: statistical approximation, functional expressivity, and functional approximation (Shumailov et al., 2024).
The first error arises because the empirical distribution $\hat{p}_{data}$ of the training data does not accurately match the true distribution $p_{data}$ [^5] The probability distributions are illustrated in Figure 6 in the appendix.. In the first iteration step, when models are trained on subsets of the original, non-model-generated training dataset, this limited representativeness is solely due to finite sampling, and the error would vanish as the sample size approaches infinity. However, in subsequent iteration steps, the training data is no longer merely a subset of the original dataset but is increasingly contaminated with examples outside the support of the true distribution, generated in previous iterations. Consequently, the representativeness of a training set depends on both recall (the fraction of the true distribution covered by the training data) and precision (the fraction of training examples that actually come from the true distribution).
The second source of error concerns the models’ capacity to collectively represent the training data distribution $\hat{p}_{data}$. For illustration, consider fitting a mixture of two Gaussians with a single Gaussian in the case of a single model, compared to fitting the same mixture with two Gaussians in the case of a collection of two diverse models (example borrowed from Shumailov et al., 2024).
The third source of error concerns the models’ capacity to learn what they are capable of representing. In our experiments, this error can be neglected because all models share the same architecture and training setup.
When all models are trained on subsets of the original training dataset, we operate under fixed precision; therefore, recall decreases as $n$ shrinks with the number of models $M$. The resulting reduction in statistical approximation for models trained on smaller datasets explains the negative effect of diversity, particularly evident in the first iteration.
At the same time, because individual models are forced to specialize in subpopulations, increasing the number of models improves functional expressivity of their ensemble. The widths of perplexity distribution reflect the collections’ expressivity, illustrated in Figure 5. Improved expressivity can increase the representativeness (recall and precision) of the collectively generated data, thereby reducing the statistical approximation error even when individual models are trained on fewer data points. Conversely, once the quality of the artificially generated training data drops because its contamination with examples outside the support of the true distribution, increasing the sample size $n$ per model can become harmful for subsequent training (compare to Sorscher et al., 2023; Xie et al., 2023; Feng et al., 2024a). This dynamic likely explains the strong performance decay of the $M=1$ ecosystem toward higher number of iterations shown in Figure 4. Thus, the improved expressivity and its downstream benefits explain the positive effect of diverse collections on model performance.
In essence, less data reduces the statistical approximation capacity of diverse collections when training data quality is sufficiently high. Simultaneously, however, diversity increases expressivity, which can improve data quality in later iterations and thereby improve statistical approximation in subsequent steps.
6. Discussion
In our study of AI ecosystem evolution, we find an optimal degree of diversity greater than zero for the performance of recursively trained models (see Figure 3). The identified optimal degree of diversity in ecosystems of models offers two key takeaways: 1) multiple community models are more successful at preserving the original information than a single model, even when trained on smaller sample sizes; however, 2) the benefits of epistemic diversity peak, but once the number of models becomes too large (and the subsets too small), this capacity decreases again. In the face of AI monoculture, our findings suggest that diversifying models is an essential factor, alongside existing data curation strategies (Xing et al., 2025; Dohmatob et al., 2024b; Feng et al., 2024a; Sorscher et al., 2023; Xie et al., 2023), for preventing collapse in human, AI-shaped knowledge.
In our experiments, conducted with small models and datasets under controlled conditions, increased diversity does mitigate, but does not entirely prevent, model collapse. Given the simplified nature of our setup, it is possible that, in more realistic scenarios, the benefits of diversity could be even more pronounced. In real-world deployments, both models and datasets are far larger, and diversity can arise from additional dimensions of system design, including model architecture, alignment approaches, user interface design, and system messaging. Combining the variations of these dimensions could amplify the benefits of diversity overall. It is also important to consider that real AI ecosystems consist of multiple but similar models. It is often assumed [^6] Of course, we wouldn’t have to assume if information about the training datasets of dominant AI models were made accessible to independent researchers. that these models are trained on the majority of accessible internet (Villalobos et al., 2024; Mozilla, 2024). If this true, then segmentation through reduction, which is the approach we took in this study, is the only approach to training-data diversification for these larg-scale, real-world AI ecosystems. If this is not the case, then diversity could be achieved through expansion by having, say, Gemini train on one portion of the internet while ChatGPT trains on another portion of the internet. Moreover, a more realistic evaluation of community models would use a community-specific, rather than a universal ground-truth, dataset, which would reduce the measured perplexity. Finally, prevailing AI systems today (e.g., OpenAI’s ChatGPT, Google’s Gemini, Anthropic’s Claude) are based on much larger models trained on significantly larger datasets. Prior work indicates a counterintuitive negative relationship between model size and stability when evaluated on a single model (Herel and Mikolov, 2024; Wright et al., 2025; Dohmatob et al., 2024a). Overall, the relationship between these factors and collapse remains insufficiently understood. Examining how the location of the optimal diversity point shifts in more realistic settings and along additional dimensions is an important direction for future research.
Our findings indicate that diversifying AI is essential for long-term knowledge production. This work contributes to a growing body of AI research that challenges the dominance of single, universal models, highlighting the benefits of diversity in contexts such as justice (Jain et al., 2024; Blili-Hamelin et al., 2025; Gebru and Torres, 2024) and safety (Lu and Kleek, 2024). Knowledge collapse is a gradual process that will likely be difficult to detect, especially in early stages. We therefore welcome recent approaches to monitoring (the progress of) epistemic diversity within models and data (Wright et al., 2025; Zhang et al., 2025a) and recommend extending such analyses to the ecosystem level. To not only monitor but also steer the ecosystem of models, a discussion that includes the public is necessary. AI development relies on public resources, not least in the form of training data. Given the high stakes, a discussion among researchers, industry, policymakers, and the public is needed to determine how these resources should be distributed to ensure a just and successful AI ecosystem in the long run.
Appendix A Probability distributions
Figure 6 illustrates the probability distributions of the information space.
Figure 6. Illustration of probability distributions of information space. Because the training data is finite, the empirical distribution, $\hat{p}_{data(t=0)}$, diverges from the true distribution, ${p}_{data}$. The model’s inherent biases (functional expressivity and functional approximation errors) further distort the distribution it represents, which is $p_{\theta(t=0)}$, and the distribution of the data it generates, which is $\hat{p}_{data(t=1)}$.
Appendix B Model progress including performance on original models
Figure 7 shows the perplexity progress of the two model types across ten iteration steps, including the models’ perplexity for the off-the-shelf models.
(a) OPT-125m
(b) GPT2
Figure 7. Comparison of ecosystems’ perplexity. In this figure, the comparison is made by increasing diversity ($D=M$) across ten iteration steps, including models’ perplexity for the off-the-shelf models (t=1t=-1). The significant perplexity drop at the first iteration step shows that the fine-tuning mechanism works.