### v2

# Epistemic diversity across language models mitigates knowledge collapse

Abstract.
As artificial intelligence (AI) becomes more widely used, concerns are growing that model collapse could lead to knowledge collapse, i.e. a degradation to a narrow and inaccurate set of ideas. Prior work has demonstrated single-model collapse, defined as performance decay in an AI model trained on its own outputs. Inspired by ecology, we ask whether increasing AI ecosystem diversity (i.e., the number of distinct models) can mitigate such collapse. To study the effect of diversity on model performance, we extend the single-model approach by segmenting the training data across an increasing number of language models and evaluating the resulting ecosystems of models over ten self-training iterations. We find that training a single model on the entire dataset improves performance only in the short term but amplifies collapse over longer horizons. Specifically, we observe that the optimal diversity level (i.e., the level that maximizes performance) increases monotonically with the number of self-training iterations. The observed effect is robust across various experimental settings, including different model families, parameter sizes, mixing human- and model-generated data, and temperature sampling methods, demonstrating the significance of ecosystem diversity for mitigating collapse. Moreover, our experiments with increased model and dataset sizes indicate that scaling up the system can amplify collapse in highly homogeneous ecosystems, thereby increasing the diversity benefits. In light of AI monoculture concerns, our results suggest the need to monitor (dis)agreement among AI systems and to incentivize more domain- and community-specific models to ensure successful knowledge production in the long run.
artificial intelligence, epistemic diversity, knowledge collapse, model collapse, AI ecosystem, system pluralism, AI pluralism
This work is supported by the University of Washington’s Center for an Informed Public and the John S. and James L. Knight Foundation (G-2019-58788).
1. Introduction
Figure 1. Illustration of model collapse in AI ecosystems. Because of a model’s (m) inherent biases (in learning and resampling) and finite data, it cannot perfectly regenerate the information space (e.g., concepts, claims, facts) it was trained on. Some information becomes distorted, and some is lost entirely due to low sampling probabilities. When this training–generation cycle is repeated, the model’s output becomes increasingly homogeneous and degrades over iterations $t$. The probability distribution of the information space expressed in the output (and represented by the model, $p_{\theta}$) gradually narrows with increasing $t$, converging toward delta-like functions. Collapse can occur at both the individual model level and the ecosystem level, meaning that even a collection of models trained on their collective output can eventually degenerate (here illustrated with three models). In society, we expect to advance knowledge. Without human intervention, the best we can hope for from AI models, though, is that they preserve the original distribution of information. Thus, for the purpose of this paper, we evaluate ecosystems based on model performance on a fixed ground-truth dataset over 10 self-training iterations.
The growing role of Artificial Intelligence (AI) in knowledge dissemination and curation raises concerns about a collapse in human knowledge (Peterson, 2025; Sourati et al., 2025; Wagner and Jiang, 2025). The scenario unfolds as follows (Figure 1): as AI systems become widely incorporated in private and professional contexts (e.g., in scientific discovery and writing (Liang et al., 2025; Hao et al., 2026; Messeri and Crockett, 2024)), they also generate an increasing volume of artificial data used for training. This amplifies the risk of model collapse, a degenerative process in which generative models re-trained on their own outputs begin producing homogeneous, biased, and nonsensical outputs (Herel and Mikolov, 2024; Shumailov et al., 2024; Bertrand et al., 2024). Ultimately, such homogenization in model-generated data, particularly with language, may contribute to knowledge collapse, a degradation to a narrow, and potentially inaccurate, set of ideas (Peterson, 2025; Wright et al., 2025; Messeri and Crockett, 2024).
Independent of these recent AI-related concerns, philosophers have argued for the benefits of epistemic diversity in science, not only for advancing knowledge production, but also to prevent injustices such as the undervaluing of minority perspectives (Douglas, 2009; Longino, 1990; Fricker, 2007; Rolin, 2019; Solomon, 2006). When scientists learn from both their own results and one another’s findings, diversity reduces the risk of converging on false beliefs. For example, monoculture thinking in medicine prolonged the myth that sickle cell disease was limited to Black Americans, leading not only to discrimination against the Black community but also to reduced treatment options for anyone affected by the disease. (Anderson et al., 2023).
“If diversity is good, how much diversity is best?” (Solomon, 2006). Simulations of collective knowledge production identify an upper bound on diversity: once diversity increases to the point where scientists no longer learn from one another, overall success rates begin to decline [^1] To clarify our position: We do not believe this upper bound has ever been reached. (Zollman, 2010; Wu and O’Connor, 2023; Fazelpour and Steel, 2022; Pöyhönen, 2017).
Like scientists exchanging new findings, AI models learn from their own output and the output of other AI [^2] We are not attempting to compare humans and AI models underlying learning mechanisms but only to illustrate information interactions in human and AI ecosystems. We use terms like “learning” and “training” because they are established in AI research but recognize the problems in anthropomorphizing this technology.. In the presence of AI monoculture—where a few large models, trained on similar datasets with similar model architectures, dominate the AI ecosystem (Lu and Kleek, 2024; Lazar and Nelson, 2023; Sourati et al., 2025; Bommasani et al., 2022; Zhang et al., 2025a)—and where evidence of model collapse exists, we ask whether increasing ecosystem diversity (i.e., the number of dissimilar models) could mitigate monocultural shifts, and whether this happens monotonically or to an optimal degree.
In a homogeneous ecosystem, collapse occurs because of a model’s non-zero probability of losing information or learning wrong patterns, which is irreversible, when relying only on its own output (Shumailov et al., 2024). Analogous to productive disagreement among humans, we hypothesize that dissimilar models trained on their collective outputs can improve collective performance by correcting each other’s errors. Viewing an AI model as an epistemological representation, we use the Hill-Shannon Diversity (HSD, $D$) from biology (Roswell et al., 2021; Jost, 2006; Hill, 1973) to measure ecosystem diversity, which quantifies the effective number of distinct AI models. It is important to note that the diversity across models serves as the primary independent variable that we manipulate to examine its effect on performance. This stands in contrast to prior work, which has focused on diversity in data (e.g., Schaffelder and Gatt, 2025; Zhu et al., 2025) and/or has treated diversity as an outcome variable (e.g., Wright et al., 2025).
Given the enormous resources required to train and operate prevailing models (e.g., ChatGPT, Gemini, Claude, etc.), studying the impact of diversity on AI ecosystem evolution at real-world scale is infeasible. And, even if we had infinite resources, many confounding factors could shape model performance, including model and data size (Kaplan et al., 2020; Bertrand et al., 2024), model architecture, and system design. To address these challenges, we build upon previous work that analyzes a simple, single-model in a semi-controlled environment (e.g. Shumailov et al., 2024; Herel and Mikolov, 2024). This previous work re-trains small language models on their own output over multiple iterations and measures performance using perplexity. We take a similar approach, but instead focus on ecosystems of models (i.e., collections of models), rather than a single model. This allows us to study the effects of ecosystem diversity on model performance, which we measure separately on two distinct model families (OPT (Zhang et al., 2022) and GPT2 (Radford et al., 2019)).
While models can be differentiated in many ways, (e.g., by size, architecture, alignment, system messaging) [^3] A specific model architecture can be used to build different AI systems (i.e., the end products with which users interact) (Brown et al., 2005). For simplicity, we use “model family” to differentiate between OPT and GPT2 pre-trained base models and “model” to refer to the fine-tuned versions., we instantiate ecosystem diversity through variation in training data.. To do this under a controlled setting, we segment the training data, which is fixed in size and, in the first iteration step, also in content. Through this process, we go from one, pre-trained model to multiple fine-tuned models (MM=1, 2, 4, 16), as shown in Figure 2. We use Wikitext2 (Merity et al., 2016) as the dataset for generating artificial training data, evaluation, and initial training. This segmentation approach allows us to vary ecosystem diversity ($D=M$) in a relatively simple manner, while keeping factors, such as model family and the per-ecosystem resource budget for sample size and computation, fixed. In our primary experimental setting, we use the smallest versions of the two model families (OPT-125m and GPT2).
Our experiments identify an optimal level of ecosystem diversity for model performance, which increases monotonically with the number of self-training iterations (Figure 3). In other words, the prevailing strategy of training a single model on as much data as possible is short-sighted: over longer horizons, a collection of models trained on segmented datasets outperforms a single model. As the number of self-training loops increases, stronger segmentation is required to reach the diversity level that maximizes overall performance. Put simply, the benefits of diversity grow with the number of iterations.
To assess the effect of ecosystem diversity under more realistic conditions, we introduce two variations of our primary setting. In the first variation (V1), we increase the size of system (i.e., model and data size). Scaling up leads to a stronger collapse in more homogeneous ecosystems, while collapse is attenuated in more diverse settings, resulting in increased diversity benefits. The second variation (V2) enhances the training data by (1) incorporating 10% fresh, original Wikitext2 data at each iteration and (2) by applying temperature sampling with $\tau=0.5,1,2$. Prior work suggests that both real-data incorporation (Shumailov et al., 2024; Gerstgrasser et al., 2024; Xing et al., 2025) and data diversification (Wright et al., 2025; Schaffelder and Gatt, 2025; Zhu et al., 2025) can mitigate collapse. While our results confirm these findings, they also show that the effect of ecosystem diversity outweighs the effects of data quality manipulation. Overall, the experimental variations suggest that the benefits of diversity for long-term performance may persist under real-world AI ecosystem conditions, where models such as ChatGPT are substantially larger, trained on far more data, can be optimized for diverse outputs, and where human-generated data is not replaced but instead accumulates alongside artificial data.
As AI models become increasingly incorporated into major institutions such as science, education, government, and economics, a collapse toward the most dominant ideas could have severe societal implications. The results of our experiments suggest that differentiation of models could be a significant factor in mitigating such collapse.
2. Background and related work
2.1. Collapse
Model collapse has been theoretically and/or empirically demonstrated for generative models in various settings (Alemohammad et al., 2023; Shumailov et al., 2024; Herel and Mikolov, 2024; Zhang et al., 2025b; Xing et al., 2025; Dohmatob et al., 2024b). The observed effects typically involve some form of performance decay; however, the field has not yet arrived at a general definition of the phenomenon (Schaeffer et al., 2025). Our experimental approach for the single model follows Shumailov et al. (2024), who first formalized model collapse and revealed it in OPT-125m. Model collapse arises primarily from three sources of error related to training data quality, the model’s capacity to represent and accurately learn a given distribution (Shumailov et al., 2024): statistical approximation error, functional expressivity error, and functional approximation error. Prior work show that increasing model and data sizes can amplify collapse (Wright et al., 2025; Herel and Mikolov, 2024; Dohmatob et al., 2024a, b; Feng et al., 2024a), thereby contrasting established scaling laws in machine learning. Model collapse is not only a threat to generative models accidentally (or inevitably) trained on artificial internet data but also when considering approaches leveraging synthetic data for efficiency reasons (Bai et al., 2022; Goyal and Mahmoud, 2024; Feng et al., 2024a). Very recent work theorizes that model collapse will result in a collapse of knowledge, defined as the continuous narrowing down of ideas, which will lead to broader epistemic problems for our society (Wright et al., 2025; Peterson, 2025; Wagner and Jiang, 2025).
2.2. Collapse mitigation
Because of the epistemic risks of collapse, scholars have developed various strategies to prevent such performance decay, or, put differently, to ensure stability. The majority of intervention strategies focus on increasing data quality. This usually involves adding fresh non–model-generated data (e.g., Bertrand et al., 2024; Gerstgrasser et al., 2024; Fu et al., 2025), filtering or correcting artificial data (e.g., Amin et al., 2025; Dohmatob et al., 2024b; Feng et al., 2024a), or diversifying training data (Zhu et al., 2025; Wright et al., 2025; Schaffelder and Gatt, 2025). While such methods have been shown to mitigate the problem, in most experimental settings they cannot entirely prevent performance decay. In addition, it is often infeasible technically and practically to constantly source with fresh data or reliably separate real from artificial data (Xing et al., 2025; Dohmatob et al., 2024a).
2.3. AI ecosystem
Our real AI ecosystem does not consist of a single self-training model but rather a collection of models trained on both their own outputs and those produced by others. Thus, more closely aligned with the work presented here are recent approaches that seek to understand model collapse at the ecosystem level—for instance, by simulating interactions among AI models via retrieval-augmented generation (RAG) (Wang et al., 2025a), or by sourcing artificial data from multiple language models for training and evaluating a single model (Schaffelder and Gatt, 2025). In contrast, our approach focuses on diversity across models as the independent variable, involving real (non-simulated) re-training of language models on their collective output.
2.4. Diversity
Epistemic diversity in humans refers to different ways of knowing. It broadly describes the phenomenon that individuals with different background assumptions, shaped by diverse lived experiences, values, and beliefs, can arrive at different interpretations when presented with the same information or data (Douglas, 2009; Longino, 1990; Solomon, 2006).
AI models generate information according to learned probability distributions, which means each model reflects specific assumptions about what counts as relevant, likely, or meaningful knowledge; in other words, a particular epistemology. Epistemic diversity in an AI ecosystem therefore captures the extent to which multiple models, shaped by distinct data sources, architectures, or objectives, may yield divergent outputs when presented with the same input. In recent AI research, epistemic diversity within and across AI models has attracted increased interest. For example, scholars have shown that diverse models collectively make better decisions (Kleinberg and Raghavan, 2021), reduce the risk of failures (Lu and Kleek, 2024), and enable fairer representations of diverse societies (Fisher et al., 2025; Varon et al., 2024; Gebru and Torres, 2024; Jain et al., 2024; Blili-Hamelin et al., 2025).
Originating in ecology, a common method to measure such diversity is the Hill–Shannon Diversity (HSD) (Roswell et al., 2021; Jost, 2006; Hill, 1973; Wright et al., 2025), which reflects the effective number of distinct, equally common elements (e.g., models, species, or categories) in a population.
Figure 2. Experimental design: manipulating ecosystem diversity and evaluating performance across self-training iterations. We examine the performance of models $m$ within collections of increasing diversity $D$ when trained on collective output across 10 self-training iterations $t$. The number of models $M$ equals the Hill-Shannon Diversity $D$, which is manipulated through segmentation of the training data. At each iteration, we concatenate and shuffle the $M$ generated datasets, and uniformly redistribute to the $M$ models for the next step. This segmentation approach keeps the data size (i.e., number of tokens) per ecosystem NN fixed, yielding a per-model sample size of $n=N/M$. The ground-truth dataset used to measure performance, the original Wikitext2 test split, is fixed across all models and iterations.
3. Approach
We study the progress of model performance in AI ecosystems containing 1, 2, 4, and 16 models, each re-trained on their collective output over 10 iterations. All models begin as identical copies of a pre-trained model but are fine-tuned on equal-sized, non-overlapping subsets of the training data [^4] This approach using non-overlapping train sets is sometimes called bootstrap aggregation (bagging) (Dietterich, 2000; Ganaie et al., 2022) and has been applied in early work studying performance of model ecosystems (e.g., Krogh and Vedelsby, 1994). This segmentation approach enables us to isolate and examine the effect of diversity across models on their performance under fixed model architecture and a fixed per-ecosystem resource budget for data size (number of tokens, $N$) and compute. We first introduce the primary experimental setting, which demonstrates a clear positive effect of ecosystem diversity on model performance in a controlled, though limited, setting. Then, in Section 3.3, we present variations of this primary setting that aim to approximate more realistic conditions and to examine how the location of the optimum shifts with factors such as model size, data size, incorporation of real data into training, and temperature sampling. We denote the following main variables: $D$ (ecosystem diversity), $M$ (number of models per ecosystem), $N$ (fixed data size per ecosystem, i.e., number of tokens), $n$ (data size per model), $T$ (number of iterations $t$, i.e., 10), and $\mu$ (model perplexity).
Each iteration step involves an evolving training dataset used to update model weights, a fixed generation dataset used to generate artificial training data for the next iteration, and a fixed evaluation dataset used to measure model perplexity (see Figure 2). The evaluation dataset is the same for all models and corresponds to the original Wikitext2 test set. The generation datasets are unique to each model but fixed across all iterations, corresponding to randomly sampled subsets of the original Wikitext2 training data.
In the first iteration step, the training data is identical to the generation dataset for a given model. Thereafter, it evolves in content because it is sourced from the outputs of the previous iteration: At each iteration, the $M$ generated datasets are concatenated, shuffled, and uniformly redistributed to the $M$ models for the next round. Throughout the ten iterations, the training and generation datasets for each model within an ecosystem are equal in size, $n=N/M$. For example, for two models ($D;M=2$), the sizes of the training and generation datasets are each half of the original size, $n=N/2$ (see Figure 2). In the case of the single-model ecosystem ($D;M=1$), our approach parallels the methods used in prior work on single-model collapse (e.g., Shumailov et al., 2024).
Translated to the real world, the segmentation approach broadly mirrors domain- or community-specific models (e.g., Feng et al., 2024b; Tanksley et al., 2025; Costanza-Chock, 2020) trained on distinct segments of internet data. While tailored to different communities that might use different inputs, the generated outputs are collectively stored in a single data pool, from which future training samples are drawn.
3.1. Ecosystem diversity D and model performance
Drawing from ecology, we use the HSD to measure ecosystem diversity, that is, diversity across models. The HSD quantifies the effective number of distinct, equally frequent models in a ecosystem (Roswell et al., 2021; Jost, 2006; Hill, 1973). Let $M$ denote the number of models, then we calculate HSD as:
(1) $D(M)=\exp\!\left(-\sum_{m=1}^{M}w_{m}\ln w_{m}\right)$
Because of equal weight across models ($w_{m}=1/M$), the diversity simplifies to the number of models in a given ecosystem (i.e. richness):
(2) $D(M)=M$
Consistent with prior work (Shumailov et al., 2024), we compute each model’s perplexity on the original Wikitext2 test split to assess model performance. Perplexity reflects a model’s ability to complete the test set text sequences (lower is better) and indicates the degree of alignment between the model’s probability distribution $p_{\theta}$ and the true distribution $p_{data}$.
In an AI ecosystem composed of $M$ different models, we calculate the combined mean perplexity $\mu$ for a given iteration step $t$ as:
(3) $\mu_{t}=\frac{1}{M}\sum_{m=1}^{M}\mu_{m,t}\text{ , }$
whereas $\mu_{m,t}$ corresponds to the mean perplexity across the text sequences of the original Wikitext2 test split.
To compare ecosystems across iterations, we calculate an aggregated performance. Let $\mu_{t}$ an ecosystem’s mean perplexity for a given iteration step $t$, then, the aggregated mean perplexity $\mu_{T}$ across $T$ iterations equals:
(4) $\mu_{T}=\frac{1}{T}\sum_{t=1}^{T}\mu_{t}$
Finally, for a fixed number of iterations $T$ and a set of ecosystems with varying diversity levels, we define the optimal diversity level as the one that achieves the highest aggregated performance:
(5) $D_{\mathrm{opt}}(T)=\arg\min_{D}\mu_{T}(D).$
3.2. Generation and training details
We use OPT (Zhang et al., 2022) and GPT2 (Radford et al., 2019), two open-accessible, small, and well-studied language model families used in prior work on collapse (e.g., Shumailov et al., 2024; Herel and Mikolov, 2024). For the primary approach, we use the smallest versions, OPT-125m and GPT2, with 125M and 124M parameters, respectively.
Analogous to prior work (Herel and Mikolov, 2024; Shumailov et al., 2024), we use Wikitext2 (Merity et al., 2016) as the original dataset. To reduce computational effort during training and generation, the primary approach uses a randomly sampled subset of two-fifths of the original size per ecosystem (N=0.8N=0.8M tokens), while keeping the original Wikitext2 validation and test splits for evaluation. Using a block size of 128 tokens, the sizes of the training–generation, validation, and test splits correspond to 7,500, 1,900, and 2,200 data points, respectively.
Our approach for generating the artificial dataset follows a method introduced in prior work (Shumailov et al., 2024). For each available token sequence of length 128 in the generation dataset, we prompt the model to generate a continuation sequence of equal length, using a five-way beam search. Because we iterate through the entire generation dataset, the model would produce the original generation dataset if it had zero error. At each iteration, each model is trained for 5 epochs with a learning rate of $1e^{-5}$ on a training sample of size $n=N/M$, using a block size of 128 tokens. The best-performing model on the validation set is preserved for subsequent iterations.
To study the effect of increasing parameter size, we use OPT-350M with 350M parameters and the full Wikitext2 dataset (2.1M tokens); see Section 3.3. The analogous experiments with GPT2-Medium (355M parameters) are reported in Appendix F. All experiments are carried out using the HuggingFace Transformers library (Wolf et al., 2019) and the Python Lightning framework (Falcon and The PyTorch Lightning team, 2019).
3.3. Variations of primary setting
To understand the effect of diversity in more realistic settings, we examine how the location of the optimum shifts under two types of experimental variations: V1 examines the effect of increasing model and data sizes, and V2 examines the effect of data quality improvements through real-data incorporation and temperature sampling. We conduct the experimental variations using the OPT model family and report the scaling experiments with GPT2 in Appendix F.
V1: System scale –– model size and data size
In reality, both model and data sizes are far larger. OPT-125m contains 125M parameters, whereas current large language models (e.g., OpenAI’s ChatGPT, Google’s Gemini, Anthropic’s Claude) likely contain in the order of trillions of parameters. A similar discrepancy applies to dataset size: while Wikitext2 contains roughly 2M tokens, the datasets used for prevailing models are likely $10^{6}$ times larger. Established scaling laws suggest a power-law relationship between performance and parameter size (Kaplan et al., 2020; Hoffmann et al., 2022). However, recent work indicates that these relationships may break down under self-training conditions and that larger models and datasets may even amplify collapse (Dohmatob et al., 2024a, b). To examine ecosystem diversity through the lens of scaling laws we introduce three variations of the primary setting: (1) a larger model (OPT-350m, with 350M parameters), (2) the full Wikitext2 training set (2.1M tokens), and (3) the combination of both larger model and dataset sizes. We conduct these experiments for M=M=1, 2, 4, 8, and 16 models.
V2: Training data quality –– real-data incorporation and sampling temperature
In realistic self-training loops, artificial data does not replace but rather accumulates alongside real (non-model-generated) data, improving the quality of the training set and mitigating collapse (Shumailov et al., 2024; Gerstgrasser et al., 2024; Xing et al., 2025). Similarly, prevailing language models are not deterministic but incorporate randomness that diversifies outputs and can mitigate model collapse (Wright et al., 2025; Schaffelder and Gatt, 2025; Zhu et al., 2025). A common mechanism is temperature sampling, which increases output diversity by flattening token-sampling probabilities (Chung et al., 2023; Alihosseini et al., 2019). To test these factors, we introduce an experimental variation that replaces 10% of the model-generated data with randomly sampled original Wikitext2 training data at each iteration (following Shumailov et al. (2024)), as well as three additional variations using sampling temperatures of $\tau=0.5$, $1.0$, and $2.0$. As a reminder, the primary experiment uses deterministic generation with $\tau=0$. When temperature sampling is applied, all remaining sampling parameters follow the Transformers default settings, corresponding to top_k = 50 and top_p = 1.0.
4. Results
We conducted a series of experiments to examine how diversity across models affects performance when models are trained on their collective outputs. Under a fixed data size and compute budget, we find that the optimal degree of diversity increases monotonically with the number of self-training iterations:
(6) $T_{1}<T_{2}\Longrightarrow D_{opt}(T_{1})\leq D_{opt}(T_{2}),$
where the optimal level is defined as the ecosystem achieving the highest aggregated performance (i.e., lowest aggregated mean perplexity). In other words, as the number of self-training iterations increases, maximizing overall performance requires segmenting the data across more models. Correspondingly, the benefits of diversity relative to the most homogeneous setting grow with the number of iterations.
In the primary approach, the single-model setting ($D=1$) performs best only in the short term (considering fewer than four iterations for OPT-125M and fewer than three iterations for GPT2, see Figure 3) Figure 4 shows the corresponding mean perplexity trajectories for ecosystems consisting of 1, 2, 4, and 16 models trained on equal-sized, non-overlapping segments of model-generated data across 10 iterations. Ecosystems with lower diversity ($D;M=1,2$) perform well initially but exhibit rapidly increasing perplexity as the number of iterations grows. In contrast, the most diverse ecosystem ($M=16$) starts with relatively high perplexity but improves over time. Across the full 10 iterations, a diversity level of four ($D;M=4$) yields the lowest aggregated mean perplexity for both model families, with a stronger effect for OPT-125M. The least diverse ecosystem ($D=1$) performs best only at $t=0$, when ecosystems differ solely in sample size per model ($n=N/M$), as all models are trained on the original data [^5] The perplexity progression, including the perplexity of the original models, is shown in Figure 8 in the appendix..
Figure 10 in the Appendix extends the analysis to 20 iterations for OPT-125M, showing that the optimal diversity level continues to increase beyond 10 iterations.
Figure 3. The optimal level of ecosystem diversity (Hill–Shannon diversity $D$) increases monotonically with the number of self-training iterations $T$. The most homogeneous setting ($D=1$, green) performs best only in the short term ($T\leq 3$ for OPT-125M and $T\leq 2 f$or GPT2). Thus, diversity $D$ mitigates model collapse. As $T$ grows, the data must be segmented across models to reach the diversity level that maximizes overall performance on the original test set ($D_{opt}(T)={\mathrm{argmin}_{D}}\,\mu_{T}(D)$). The corresponding performance trajectories are presented in Figure 4. Colored symbols correspond to the diversity levels shown in Figure 2 1, 2, 4, and 16 (e.g., orange denotes $D=2$). Intermediate levels are omitted, as the key result is the monotonic increase in diversity benefits with $T$. An extended range of up to 20 iterations is shown in Figure 10 in the Appendix.
(a) OPT-125m
(b) GPT2
(c) Aggregated perplexity after 10 iterations
Figure 4. Ecosystem diversity mitigates collapse. (a, b) Comparison of ecosystems’ mean perplexity with increasing diversity ($D=M$) at fixed dataset size across ten iteration steps. While low-diversity ecosystems ($D=1,2$) exhibit strong performance decay, the high-diversity ecosystem ($D=16$) starts with relatively high perplexity but improves over $t$ . The shaded region is a combined standard deviation across the models and data points. (c) Aggregated perplexity. Both model families exhibit an optimal level of diversity of 4 for overall performance across 10 iterations. Colors correspond to the diversity levels shown in Figure 2 (e.g., orange denotes $M;D=2$).
(a) Optimal diversity levels.
(b) Aggregated perplexities after 10 iterations.
(c) Mean perplexities for M;D=1
(d) Mean perplexities for M;D=16
[…]
(a) Optimal diversity levels.
(b) M;D=1
(c) M;D=4
[…]
(a) Optimal diversity levels.
(b) M;D=1
(c) M;D=16
[…]
4.1. V1: Model size and data size
The benefits of diversity increase as both model size (from 125M to 350M parameters) and dataset size (from 0.8M to 2.1M tokens) scale (Figure 5). Holding dataset size constant, larger models exhibit a steeper increase in the optimal diversity level as the number of iterations grows. A similar pattern appears when increasing dataset size for the larger model, whereas for the smaller model dataset size has only a minor effect on the optimal diversity level. Below, we highlight several notable performance effects relative to the primary setting (125M-parameter model and 0.8M tokens dataset).
At iteration step $t=0$, when models are trained on the original dataset, increasing either model or dataset size improves performance in line with established scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022). The most pronounced effects of larger models is the amplified collapse of low-diversity ecosystems ($D\leq 2$), leading to the observed shift in the optimal diversity level. While the effects of dataset size are less pronounced, two patterns emerge: increasing dataset size for the the larger model improves performance primarily in the most homogeneous ecosystem ($D=1$), whereas for the smaller model it improves performance primarily in the most diverse setting ($D=16$).
4.2. V2: Real-data incorporation and sampling temperature
Improving data quality by incorporating 10% real data (Figure 6) or by applying temperature sampling with $\tau=1.0$ and $\tau=2.0$ (Figure 7) generally improves model performance. Because both interventions primarily benefit the most homogeneous ecosystem ($D=1$), performance differences across diversity levels shrink relative to the primary setting. However, since the positive effect of ecosystem diversity exceeds these improvements, both interventions produce a similar trajectory of the optimal diversity level over iterations as observed in the primary setting. We highlight several notable performance effects relative to the primary setting below.
The impact of incorporating real data at each iteration increases as diversity decreases, reflecting the larger per-model dataset size ($n=N/M$). Temperature sampling can also affect data quality by diversifying generated outputs. For temperatures of $\tau=1.0$ and $\tau=2.0$ in the most homogeneous ecosystem ($D=1$), perplexity increases during early iterations but declines by a similar magnitude later in the loop, resulting in only marginal overall improvement. In more diverse ecosystems, $\tau=1.0$ produces slightly negative effects, while $\tau=2.0$ leads to strongly negative effects. By contrast, a temperature of $\tau=0.5$ shows almost no effect relative to the primary setting for any fixed diversity level.
5. Theoretical interpretation
Across experiments, our study of ecosystem evolution reveals a monotonic increase of the optimal diversity level with number of self-training iterations $T$. For a fixed number of 10 iterations, we find a U-shaped relationship between diversity and collapse. Considering that dataset size per model $n=N/M$ decreases as diversity increases, the observed increase in optimum diversity level indicates a complex interplay between opposing effects of diversity and scale with $T$. The complexity grows when considering the results obtained with increased scale (Section 4.1): When model size increases, we observe stronger collapse in low-diversity ecosystems ($D\leq 2$) and weaker collapse in high-diversity ecosystems ($D\geq 4$). Increasing dataset size generally improves performance for OPT-125m but worsens performance for OPT-350m, except in the $D=1$ ecosystem. In this section we propose a theoretical explanation of the underlying effects.
Dohmatob et al. (2024a) theorize that increasing model and/or dataset size amplifies model collapse once data quality becomes sufficiently low. Building on this theory, we can assume that low-diversity ecosystems ($D\leq 2$) tend to operate in a low data-quality regime, leading to stronger collapse as model size increases, while the opposite holds for high-diversity ecosystems ($D\geq 4$). This raises the question of how diversity helps preserve high data quality, given that diverse ecosystems start with lower model performance at iteration step $t=0$. To address this, we introduce effective data quality (EDQ), which builds on the framework of Dohmatob et al. (2024a). We then use this hypothesis to explain our empirical results in detail.
Effective data quality EDQ
Training data quality reflects the discrepancy between $\hat{p}{data}$ and pdatap{data} (Dohmatob et al., 2024a; Wang et al., 2025b), for example as measured by (Shumailov et al., 2024; Bertrand et al., 2024): $\delta(\hat{p}_{data})=\mathbb{W}_{2}(\hat{p}_{data},p_{data})$, whereas $\mathbb{W}_{2}$ refers to the Wasserstein-2 distance between the two distributions [^6] The probability distributions are illustrated in Figure 11 in the appendix.. For an original (non–model-generated) dataset, all examples lie within the support of the true distribution, and this discrepancy approaches zero as dataset size approaches infinity. In contrast, due to models’ internal errors (functional expressivity and functional approximation errors (Shumailov et al., 2024)), resampled artificial datasets can contain less information about the true distribution and can even include examples outside its support. Moreover, depending on the data received in previous training steps, not every in-distribution example provides new or useful information to a model (Feng et al., 2024a; Sorscher et al., 2023) and further training on a dataset similar to the initial training set risks performance decay via overfitting (Tirumala et al., 2022). Effective fine-tuning methods suggest that optimal training data is not merely data that reflects the true distribution, but data that most efficiently moves model weights toward that distribution (Ouyang et al., 2022; Liu et al., 2025; Ji et al., 2024; Deb et al., 2025; Antonello et al., 2021). Hence, optimal data must account for the model’s existing distribution, which approximates a mixture of previous training distributions, i.e., ${p}_{\theta(t)}\approx\sum_{i=0}^{t-1}\hat{p}_{data(i)}$. We define EDQ at iteration $t$ as the difference between the training–truth discrepancy and the model–truth discrepancy [^7] Because we introduce EDQ primarily to explain our empirical results, we rely on a simple scalar metric to capture distributional discrepancy and ignore additional factors required for optimal model weights updates, such as the direction of the distributional shift.:
(7) $\delta_{EDQ}(\hat{p}_{data},{p}_{\theta)}:=\mathbb{W}_{2}(\hat{p}_{data(t)},p_{data})-\mathbb{W}_{2}({p}_{\theta(t)},p_{data}),$
where lower values of $\delta_{EDQ}$ indicate higher EDQ. EDQ reflects the model-dependent informativeness of the given training data. Using $\delta_{EDQ}$, we can now state our main hypothesis under finite training data and compute:
Hypothesis: EDQ, informal
  1. (1)
    High-EDQ regime: If $\mathbb{W}_{2}(\hat{p}_{data(t)},p_{data}))$ sufficiently lower than $\mathbb{W}_{2}({p}_{\theta(t)},p_{data})$, additional training improves performance.
  2. (2)
    Low-EDQ regime : If $\mathbb{W}_{2}(\hat{p}_{data(t)},p_{data}))$ sufficiently greater than $\mathbb{W}_{2}({p}_{\theta(t)},p_{data})$, additional training degrades performance.
  3. (3)
    Medium-EDQ regime: If $\mathbb{W}_{2}(\hat{p}_{data(t)},p_{data}))\approx\mathbb{W}_{2}({p}_{\theta(t)},p_{data})$, additional training neither significantly improves nor degrades model performance.
This hypothesis has two implications for recursively trained models. First, performance gains or losses depend on EDQ, regardless of whether the data is generated by humans or models (Wang et al., 2025b; Feng et al., 2024a), or on its diversity per se. Nevertheless, both the proportion of artificial data and data diversity can shape the EDQ. Second, when operating in a low-EDQ regime, neural scaling laws invert: performance decay increases with system scale, consistent with prior findings (Dohmatob et al., 2024a; Herel and Mikolov, 2024; Wright et al., 2025). Neural Scaling laws state that test loss depends on model size and dataset size, following a power-law relationship (Kaplan et al., 2020; Hoffmann et al., 2022).
Diversity and EDQ
The EDQ hypothesis helps explain the impact of ecosystem diversity on model performance. First, in our segmentation approach, sample size per model decreases as diversity increases; thus, once models operate in a low-EDQ regime, greater diversity becomes beneficial. This relationship is consistent with prior work (Dohmatob et al., 2024a) and reflects a trade-off in long-term considerations of AI ecosystems between large-scale under high-EDQ and small-scale under low-EDQ. Second, the distribution of an artificial dataset closely resembles the distribution of the model from which it is sampled. If a model had zero sampling error, these distributions would be identical. However, under our hypothesis, data whose distribution closely mirrors the model itself has low EDQ. In contrast, in a diverse ecosystem, different models represent different components of the true distribution. As a result, the EDQ of redistributed outputs can be substantially higher than in a setting where models only see their own output: data that is not useful for the model that generated it may still be useful for another model within the same ecosystem (see Illustration 12 in the Appendix). This introduces a second trade-off for maximizing EDQ: between larger datasets per model when the data are fresh to the model ($t=0$) and higher EDQ through diversity in later iterations.
We now use the proposed theory to interpret performance dynamics across different stages of self-training iterations.
Iteration step $t=0$ At $t=0$, all models are trained on subsets of real data and therefore operate in a high-EDQ regime. EDQ increases with sample size per model $n=N/M$, which decreases as the number of models $M$ increases. Because default scaling laws apply in this regime, performance decreases with diversity and improves with system scale (see Section 4.1).
Later iterations In later iterations, more diverse ecosystems can exhibit lower performance loss (at $t=1$) and/or greater performance gains because (i) the exchange of artificial data among models increases EDQ, and (ii) once models enter the low-EDQ regime, larger per-model sample sizes $n$ become harmful. This explains the sharp decay observed for $D=1$ in Figure 4, which is further amplified by larger models (Figure 5). Performance trajectories after $t=0$ illustrate that models in diverse ecosystems receive data of higher EDQ than models in homogeneous ecosystems, ultimately yielding superior long-term performance, despite lower model performance at the beginning.
Post hoc sensitivity analyses of individual factors that may influence the performance dynamics observed in the evolution of AI ecosystems support our EDQ hypothesis (see Appendix H). The results indicate that neither (model-independent) data quality, training data size, the number of models per ecosystem, nor overfitting can fully account for the observed performance decay or the differences across ecosystems.
6. Discussion
In our study of AI ecosystem evolution, we find that the optimal level of diversity increases monotonically with the number of self-training iterations (see Figure 3).
This offers a new perspective on generative model performance. Training a single model on as much data as possible is short-sighted: over longer horizons, community models trained on segments of the available data outperform a single model with access to the full dataset. The more often artificial data is recycled for training, the greater the degree of segmentation needed to attain the optimal diversity level. Put simply, the benefits of diversity grow with the number of iterations.
This observed effect is robust across model families, model sizes, dataset scales, and experimental variations involving real data and temperature-based sampling. Moreover, we find that the benefits of diversity also grow as both models and datasets scale. Taken together, our results suggest that differences between models are a key mechanism for mitigating collapse even under more realistic conditions where models are larger, trained on far more data, can be optimized for diverse outputs, and where human-generated data accumulates alongside model-generated data.
Before turning to sociotechnical implications for AI evaluation and governance, we briefly restate the technical aspects from Section 5. To interpret our empirical results, we introduce effective data quality (EDQ), which treats the informativeness of data as a function of both the model distribution and the true distribution [^8] Importantly, even if the EDQ hypothesis were falsified, the empirically observed diversity optimum would remain.. When EDQ is sufficiently high, additional training improves performance; when EDQ is low, further training is harmful and performance declines. This framing has two broad implications for recursively trained models and machine learning more generally. First, gains and losses depend on EDQ regardless of whether the data is human- or model-generated (Wang et al., 2025b; Feng et al., 2024a; Antonello et al., 2021). Second, in low-EDQ regimes, standard scaling laws invert: performance degradation increases with scale, consistent with prior findings (Dohmatob et al., 2024a; Herel and Mikolov, 2024; Wright et al., 2025). Through the lens of EDQ, diversity helps prevent collapse for two reasons: (i) once the low-EDQ regime is reached, reducing data size per model via segmentation becomes beneficial; and (ii) distributing model-generated data across diverse models increases the model-dependent informativeness of the data each model receives. By contrast, increasing data per model (and reducing diversity) is only beneficial early on, when models are first exposed to new data and EDQ is dominated by sample size. The advantage of monoculture observed short-term can even disappear entirely under three modeling variations: introducing diversity by combining model families rather than segmenting data, instantiating diversity via dataset expansion rather than dataset segmentation, and evaluating community models on community-specific test sets rather than a universal benchmark. These findings, reported in the Appendix G, suggest that multiple forms of model diversity potentially can potentially be combined to mitigate collapse.
How should these results inform AI goals and governance? Most directly, the long-term epistemic risks of AI monoculture challenge the premise of a single universally optimal model, aligning with prior calls for diversity motivated by justice, fairness, and safety (Jain et al., 2024; Blili-Hamelin et al., 2025; Gebru and Torres, 2024; Lu and Kleek, 2024; Fisher et al., 2025; Lazar and Nelson, 2023). Importantly, promoting diversity in AI does not imply endorsing inferior models; rather, it acknowledges that there are multiple legitimate conceptions of what constitutes beneficial AI.
Concerns about AI-driven knowledge collapse (Peterson, 2025; Wright et al., 2025) reflect the need of evaluation frameworks that capture long-term ecosystem effects—effects that short-term benchmark performance may miss. Our results suggest that model similarity (i.e., ecosystem homogeneity) can serve as an indicator of collapse risk. We welcome recent efforts to measure epistemic diversity both within (Wright et al., 2025) and among (Zhang et al., 2025a) models, and encourage AI operators to assess their models’ disagreement relative to other models deployed in the same ecosystem.
Policy approaches from other domains may offer useful analogies, including biodiversity protection, DEI initiatives in science, and regulations that limit excessive genetic homogeneity (e.g., cousin-marriage laws). In AI, governance could reward model outputs increasing diversity in shared online data pools or regulate the distribution of critical resources such as training data. Diversity can also be incentivized bottom-up through model specificity: requiring systems to be explicit about the tasks, communities, languages, values, etc. they serve (Blili-Hamelin et al., 2025) would naturally push ecosystems toward greater differentiation. As long as models remain beneficial within their specified contexts, too much diversity should not be a concern.
Given the current trajectory toward AI monoculture (Lazar and Nelson, 2023; Sourati et al., 2025; Zhang et al., 2025a), a shift toward greater model diversity appears necessary to sustain long-term knowledge production.
Appendix A Model progress including performance on base models
Figure 8 shows the perplexity progress of the two model families across ten iteration steps, including the models’ perplexity for the base models.
(a) OPT-125m
(b) GPT2
Figure 8. Comparison of ecosystems’ perplexity. In this figure, the comparison is made by increasing diversity ($D=M$) across ten iteration steps, including models’ perplexity for the off-the-shelf models (t=1t=-1). The significant perplexity drop at the first iteration step shows that the fine-tuning mechanism works.
Appendix B OPT-125m M=1 runs with different random seeds
Figure 9 shows the performance trajectories for $M=1$ across ten iterations for four independent runs. The similarity of these trajectories indicates that the results are largely independent of the random seed. Ecosystems with higher diversity already average performance across two or more models.
Figure 9. The performance trajectory of $M=1$ is largely run-independent. Mean perplexity and standard deviation for ecosystem $M=1$ for each of four independent runs with different random seeds.
Appendix C OPT-125m performance across 20 iterations
Figure 10 shows the optimal diversity level for OPT-125m across up to 20 self-training iterations ($T$), showing that the optimal diversity level continues to increase beyond 10 iterations.
(a) Optimal level of diversity.
(b) Performance progress.
Figure 10. The optimal level of ecosystem diversity continues to increase beyond 10 self-training iterations ($T$). (a) Optimal diversity level and the corresponding performance trajectories for OPT-125M. (a) and (b) extend the iteration range of Figures 3 and 4. Colored symbols correspond to the diversity levels shown in Figure 2 with an additional level of $D=8$ in light green.
Appendix D Illustrations of probability distributions and EDQ
Figure 11 illustrates the probability distributions in self-training loops, and Figure 12 illustrates the model-dependent EDQ.
Figure 11. Illustration of probability distributions in self-training loops. Because the training data is finite, the empirical distribution, $\hat{p}_{data(t=0)}$, diverges from the true distribution, ${p}_{data}$. The model’s inherent biases (functional expressivity and functional approximation errors) further distort the distribution it represents, which is $p_{\theta(t=0)}$, and the distribution of the data it generates, which is $\hat{p}_{data(t=1)}$.
Figure 12. Illustration of model-dependent EDQ. For the illustrated distribution discrepancies, the training data (“DATA”) is beneficial for “m2” (high EDQ) but harmful for “m1” (low EDQ). EDQ is defined as $\delta_{EDQ}(\hat{p}{data},p{\theta}):=\mathbb{W}2(\hat{p}{data(t)},p_{data})-\mathbb{W}2(p{\theta(t)},p_{data}),$ where lower values of $\delta_{EDQ}$ indicate higher EDQ. It is important to note that the introduced scalar metric serves as a heuristic to explain model-dependent data quality; therefore, we ignore additional factors shaping model weights.
Appendix E Perplexity distributions for both model families
Figure 13 shows the mixed perplexity distributions of the OPT-125m and GPT2 ecosystems on the original training dataset, indicating that model collections with greater diversity exhibit both greater variety in the information represented and lower alignment with the training data distribution.
(a) OPT-125m
(b) GPT2
Figure 13. Perplexity distributions widen with increasing diversity. Mixed perplexity distribution on the original Wikitext2 train data points, evaluated by all models in a given ecosystem with diversity ($D$ =$M$ = 1, 2, 4, 16) at iteration steps $t$ = 0 (left) and $t$ = 9 (right). As number of models increases, the distributions become wider, indicating greater diversity. Simultaneously, their means shift toward higher perplexity, suggesting greater divergence from the distribution in the original training set. Across all ecosystems, the distributions narrow from the beginning to the end of the model iterations.
Appendix F GPT2 V1: Model size and data size
Figure 14 shows GPT2 performance results for increased model size (GPT2-Medium) and larger data size (2.12.1M tokens).
(a) Optimal diversity levels.
(b) Aggregated perplexities after 10 iterations.
(c) Mean perplexities for M;D=1
(d) Mean perplexities for M;D=16
[…]
Appendix G Additional experimental variations of the primary setting
The observed short-term benefit of homogeneous ecosystems is due to the segmentation approach, which reduces the data size per model as diversity increases, and to the use of a universal ground-truth dataset (the Wikitext2 test set in our case). This initial advantage disappears when diversity is instantiated through combining two model families (Figure 16), through data expansion (Figure 17(b)), or when community-specific test sets are employed. The aggregated results are shown in Figure 15.
Diversity through combining 2 model families
Real-world diversity may arise from multiple system-design dimensions, including pre-training data, architecture, alignment, interface design, and system messaging. Figure 16 shows that combining two model families (OPT-125m and GPT2) consistently outperforms the combination of two models of same family, suggesting that diversity across design dimensions may further amplify benefits.
Diversity via data expansion
The expansion approach reflects a more realistic landscape of multiple similar models (Lazar and Nelson, 2023; Zhang et al., 2025a). If dominant models are trained on largely overlapping internet-scale data (Villalobos et al., 2024; Mozilla, 2024), segmentation through reduction may be the only feasible diversification strategy; otherwise, diversity could be achieved via expansion by incorporating models trained on previously unused data. In such comparisons, diverse ecosystems consistently achieve lower perplexity across iterations (Figure 17(b)), whereas simply adding more similar models does not prevent collapse (Figure 20(a)).
Community-specific test sets
Evaluating community models using community-specific datasets rather than universal benchmarks yields a more realistic assessment; under this setting, diverse ecosystems outperform homogeneous ones already at $t=0$ (Figure 17(a)). Community- or domain-specific models challenge the prevailing assumption that AI systems should serve everyone equally across all tasks (Blili-Hamelin et al., 2025; Bender et al., 2021). Granting affected communities greater agency in AI design (Costanza-Chock, 2020) may naturally guide ecosystems toward healthier diversity.
Human language and knowledge evolve alongside AI systems. Following prior work (e.g., Herel and Mikolov, 2024; Shumailov et al., 2024; Alemohammad et al., 2023), we use a static dataset for generation and evaluation; in practice, this would resemble evaluating future models on today’s data. A more realistic approach would involve evolving generation and evaluation datasets over time.
While further research is needed to identify optimal ecosystem parameters under realistic conditions, our results provide strong evidence that diversity is a key factor in sustaining long-term AI ecosystem performance.
Figure 15. The advantage of homogeneous ecosystems observed short-term can even disappear entirely under three modeling variations: introducing diversity by combining model families rather than segmenting data (“OPT-125m(+GPT2)” and “GPT2(+OPT-125m)”) instantiating diversity via dataset expansion rather than dataset segmentation (“OPT-125m, Expansion”), and evaluating community models on community-specific test sets rather than a universal benchmark (“OPT-125m, Community”). These variations involved only two diversity levels ($D=1,2$).
(a) Combination of two OPT-125m models compared with an OPT-125m paired with GPT-2
(b) Combination of two GPT-2 models compared with an GPT-2 paired with OPT-125m
Figure 16. Two diverse pre-trained models yield lower perplexity than two identical pre-trained models. OPT-125m+OPT-125m and GPT2+GPT2 represent ecosystems composed of two identical models (green). OPT-125m(+GPT2) and GPT2(+OPT-125m) represent ecosystems where each model is paired with the other type (orange). Unlike the primary segmentation approach, diversity here is instantiated by using two different pre-trained models. All models are trained on the same dataset at iteration step $t=0$ and use the same generation dataset throughout; thus, aside from their pre-training differences, the training procedure is identical across models. The shaded region represents the combined standard deviation across models and data points.
(a) From the first iteration step ($t=0$) onward, two community models consistently achieve lower perplexity than a single model trained and evaluated on the combined datasets. $M=1$: a single model trained and evaluated on the combination of Wikitext2 and Penn Treebank (green). $M=2$: two community models, each trained and evaluated on its corresponding dataset (orange). To ensure a controlled comparison, the length of the Wikitext2 dataset is reduced to match the length of Penn Treebank.
(b) Two models trained on two different segments of Wikitext2 consistently achieve lower perplexity than two models both trained on one of the two segments. $D=1$: Two models trained on identical segment 0 (green). $D=2$: Two models trained on two different segments 0 and 1 (orange), corresponding to the same ecosystem as M=2 in the primary setting. Unlike in the primary segmentation approach, diversity is achieved through expansion of the distribution.
Figure 17. The shaded region is a combined standard deviation across the models and data points.
Appendix H Sensitivity analysis
We conducted additional experiments to isolate factors that might affect the performance changes observed in the primary setting. The results indicate that neither (model-independent) data quality (Figure 18), training data size (Figure 19), the number of models per ecosystem (Figure 20), nor overfitting (Figure 21) can fully account for the observed performance decay or the differences across ecosystems. However, overfitting, which is captured by the EDQ, appears to be an important contributor to performance decay.
Because these individual factors fail to explain the performance trajectories in isolation, the findings provide support for the EDQ hypothesis.
(a) OPT-125m separately trained on generated output
(b) OPT-125m tested on generated output
Figure 18. Data quality alone cannot explain performance differences across ecosystems. The OPT-125m base model was separately trained (a) and tested (b) on generated outputs from the least and most diverse ecosystems ($D=1$ and $D=16$, respectively). For $D=16$, outputs from all 16 models were aggregated. In (a), the off-the-shelf OPT-125m base model was fine-tuned separately on each generated dataset and subsequently evaluated on the fixed Wikitext2 test set. The similar perplexity values obtained for $D=1$ and $D=16$ indicate that model-independent data quality cannot fully account for the performance differences observed in our main experiments. Lower perplexity values for outputs from iterations $t=0$ and $t=1$ in (b) suggest that artificial data generated at early stages of the self-training loop more closely resembles the pretraining distribution of the base OPT-125m model. Fine-tuning on such artificial data shifts the distribution of the fine-tuned model back toward the pretraining distribution and away from the true distribution, which helps explain the performance drop observed with the output from iteration $t=0$ in (a) and consistently across our experiments.
(a) Single OPT-125m re-trained on the real Wikitext2 train set
(b) Single OPT-125m self-trained on own output
Figure 19. Training data size alone cannot explain performance differences across ecosystems. A single OPT-125m model was analyzed in two conditions: (a) re-trained on the real Wikitext2 training set and (b) self-trained on its own output, respectively. The initial training data content and sizes correspond to the primary setting (N=0.8N=0.8M tokens), but only one model was used in each run. The condition “$D=1$, OPT-125m, NN” matches $D=1$ in the primary setting. The similar perplexity values observed across different data sizes indicate that training data size alone cannot account for the performance differences across ecosystems.
(a) OPT-125m
(b) GPT-2
Figure 20. Number of models cannot fully explain the performance differences observed across ecosystems. 1×OPT-125m and 1×GPT-2: a single model trained on the entire dataset of size $N$, corresponding to the $D=1$ condition in the primary setting (green). 2×OPT-125m and 2×GPT-2: two identical models, each trained on the entire dataset of size NN (orange). The generation datasets are identical for all models. The training datasets are identical at the first iteration step but may differ slightly in later iterations for the two-model runs due to stochasticity in setup. The shaded region represents the combined standard deviation across models and data points.
(a) OPT-125m
(b) GPT-2
(c) OPT-125m evaluated on train set
(d) GPT-2 evaluated on train set
Figure 21. Overfitting alone cannot explain the observed performance decay. Comparison of the self-training loop in the primary setting (green) with a re-training loop on the fixed real training set (orange), evaluated on the test set (top) and training set (bottom). Because the training data in the re-training condition remains fixed, any performance decay can be fully attributed to overfitting. The lower perplexity on the training set in the re-training condition indicates stronger (over)fitting to the fixed training data. This suggests that overfitting alone cannot account for the performance decline observed in the self-training loop, which appears comparable for OPT-125m and even stronger for GPT-2.
Appendix I Hyperparameters
Table 1 provides the hyperparameters used for text generation.
Hyperparameter Value
temperature 0.0
do_sample False
max_new_tokens 128
min_new_tokens 128
repetition_penalty 0.0
batch size 32
block size 128
num_beams 5
Table 1. Generation hyperparameters.